A wave traveling 77 m/s has a wavelength of 4.0 m. What is the frequency of the wave (in Hz)?
1
Expert's answer
2019-04-29T09:47:07-0400

The formula for frequency is

"f=\frac{v}{λ}"

"f=\frac{77\frac{m}{s}}{4.0m}=19 s^{-1}= 19 Hz"